A man is 18 years older than his son, in 2 years the man will be double his age
Answer by TimothyLamb(4379)

m = s + 18
m + 2 = 2(s + 2)
---
m + 2 = 2s + 4
16 = s
---
s = 16
m = 34
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php